OpenStudy (anonymous):

What is f(5) for f(x) = 4x2 – 5x + 2? 127 –23 77 150

OpenStudy (anonymous):

\[f(5)=4(5)^2-5(5)+2\]

OpenStudy (anonymous):

\(\huge f(5)=4(25)-25+2\) \(\huge f(5)=100-23\)

OpenStudy (anonymous):

So, what is \(\color{maroon}{\huge 100-23=?}\)

OpenStudy (anonymous):

77

OpenStudy (anonymous):

correcto

OpenStudy (anonymous):

\(\huge \color{pink}{\heartsuit}\)